an automobile radiator contains 16 liters of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should be drained and replaced with pure antifreeze so that there will be 50% antifreeze?
Answer by lwsshak3(11628)

an automobile radiator contains 16 liters of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should be drained and replaced with pure antifreeze so that there will be 50% antifreeze?
**
let x= amt of 30% antifreeze to be drained and replaced with pure antifreeze
16-x= amt of 30% antifreeze remaining in radiator
30%(16-x)+100%x=50%*16
4.8-.3x+x=8
.7x=3.2
x=3.2/.7
x≈4.57
amt of 30% antifreeze to be drained and replaced with pure antifreeze=4.57 liters